If a car is traveling at a speed of 28.0 m/s, is it exceeding the speed limit of 55.0 mi/h ?
1
Expert's answer
2020-09-22T15:36:04-0400

Converting "28m\/s" to "mi\/h", obtain:


"28\dfrac{m}{s} = 28\cdot \dfrac{0.000621mi}{\frac{1}{3600}h} = 28\cdot 2.24\dfrac{mi}{h} \approx 62.72\dfrac{mi}{h}"

As far as "62.72>55", then the car is exceeding the speed limit.


Answer. Yes.